OpenStudy (anonymous):

"Determine an equation of a quadratic function with x-intercepts of 2 and 6, that passes through the point (3, 12)"

OpenStudy (anonymous):

I know how to do this question if I'm given the points of the vertex.. but in this case, it isn't. so..

OpenStudy (anonymous):

start with \[f(x)=a(x-2)(x-6)\] and then solve for \(a\) via \[f(3)=a(3-2)(3-6)=12\]

OpenStudy (anonymous):

you get \[a\times 1\times -3=12\] \[-3a=12\] \[a=-4\] so your quadratic is \[-4(x-2)(x-6)\] multiply out if you want

OpenStudy (anonymous):

i get \(f(x)=-4 x^2+32 x-48\) but check my algebra

OpenStudy (anonymous):

@satellite73, I got the same thing but the answer says it should be y = - 4 ( x - 4 )^2 + 16.

OpenStudy (anonymous):

fine. i guess they wanted it in "vertex" form

OpenStudy (anonymous):

Oh yeah, nevermind i got it. thank you so much!

OpenStudy (anonymous):

convince yourself that they are equal

OpenStudy (anonymous):

yw